OpenStudy (anonymous):

ILL GIVE MEDAL AND A FAN!!! PLZ HELP ME A collection of dimes and nickels is worth $9.50. There are 132 coins in all. How many of each coin are there?

OpenStudy (anonymous):

@AnswerMyQuestions please help

OpenStudy (anonymous):

@AnswerMyQuestions you still there?

OpenStudy (anonymous):

Not so sure...

OpenStudy (anonymous):

this is so confusing to me... to me it seems like there is no way of knowing... but there might be...

OpenStudy (anonymous):

Ikr?

OpenStudy (anonymous):

@e.mccormick @myininaya

OpenStudy (anonymous):

yeesh this is hard. im looking thru my math books, google, i just cant find it

myininaya (myininaya):

ok dimes are .10 nickels are .05 we want to know how many .10+how many .05 equals 9.5 so let's translate this x(.10)+y(.05)=9.5 where x represents the number of dimes and y represents the number of nickels we are also given that the amount of coins we have is 132 so that means we also have x+y=132 since x is the number of dimes and y is the number of nickels

OpenStudy (anonymous):

ok, i think im with you so far...

myininaya (myininaya):

so try to solve the system: .1*x+.05*y=9.5 x+y=132

OpenStudy (anonymous):

oh boy, lol ok gimmie a minute

myininaya (myininaya):

if you don't like that first equation, I recommend getting rid of the decimals by multiply both sides by 100

OpenStudy (anonymous):

ok,

myininaya (myininaya):

you can write here step by step what you are doing before you move on

myininaya (myininaya):

you know so you don't do a lot of pointless work

OpenStudy (anonymous):

ok, hold on, im doing the math in my head, its just taking a while...

OpenStudy (anonymous):

so r there 60 dimes?

OpenStudy (anonymous):

wait, 61?

myininaya (myininaya):

I got something close to that. But maybe I did something wrong.

myininaya (myininaya):

yeah try writing your work here so i can see what went wrong

OpenStudy (anonymous):

no wait 59,

OpenStudy (anonymous):

im doing it all in my head. :D hang on i think i got it, if not, ill start writing it

OpenStudy (anonymous):

omg im so close

OpenStudy (anonymous):

I GOT IT!!!!!

OpenStudy (anonymous):

58 DIMES AND 74 NICKELS!!!!

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

58 x .10 = 5.8 74 x .05 = 3.7 5.8 + 3.7 = 9.5!!!

OpenStudy (anonymous):

THANK YOU SO MUCH

myininaya (myininaya):

no problem